# DriveSOTIF: Advancing Perception SOTIF Through Multimodal Large Language Models

### v1

@@ VI-B Real World Data Evaluation
For real-world evaluation, as shown in Figure VI-B, we assessed the fine-tuned models using data collected in Waterloo, Ontario, Canada, along with proprietary dash-cam footage from China. This evaluation is conducted through a ROS code that queries the MLLM API with random questions from a list, along with image data from collected data in the form of ROSBag or video.
@@
In detail, we used image and video data from the University of Waterloo’s WATonoBus project [47] and the Canadian Adverse Driving Conditions (CADC) dataset [48]. The WATonoBus is an autonomous shuttle operating along a 2.7-kilometer, two-lane University Ring Road that encircles the campus [14]. The CADC dataset was gathered by an autonomous vehicle platform navigating various areas around Waterloo.
@@
TABLE IX: Sample labeled Questions & Answers and image captions from SOTIF-VQA dataset.
@@
A: The risks include potential misidentification of road boundaries due to snow, reduced contrast between the road and its surroundings, and the presence of workers and equipment which could be obscured by snow or not well-illuminated. A: The perception-related SOTIF risk in this image is caused by heavy snowfall, which can reduce visibility and obscure road markings, signs, and other vehicles.
@@
VI-C Results and Discussion
As can be seen in Table VI-B, in our experiments, fine-tuned MLLMs performed effectively in challenging environments, yet it also highlighted critical areas for enhancement. Notably, in Waterloo’s snowy conditions, the MLLM successfully managed complex scenarios with obscured road markers and temporary obstacles like snowplows and maintenance crews. This performance highlights the model’s robust feature recognition capabilities under adverse weather conditions, proving its potential in dynamic real-world applications.
@@ VI-C Results and Discussion
This demonstrates the model’s capability to not only detect risks but also understand the causal links between challenging conditions and safety threats, providing actions that could effectively reduce risks in real-world driving situations. With only 8 GB of GPU memory, the fine-tuned Qwen2-VL 2B model can be deployed to vehicles and replace the API-based vision model for local, (near) real-time visual risk assessment. This eliminates privacy concerns, cybersecurity risks, and reliance on internet access.
Pairing this model with a decision-making framework like DriveLLM[14] enables autonomous systems to factor SOTIF risks into real-time decision-making. This integration improves safety in complex environments and potentially prevents fatal accidents like the child-hidden-in-the-box scenario, where conventional perception systems may fail to anticipate hidden risks. Most importantly, deploying a fine-tuned model that supports continuous improvement significantly enhances adaptability. By continuously updating the dataset and applying techniques like SFT and Kahneman-Tversky Optimization (KTO)[49], the MLLM can progressively refine its performance and extend to other tasks such as scenario understanding and ODD assessment.
Conversely, the MLLM’s performance in detecting subtle hazards, such as the partially visible child in a box during the China experiments, revealed some limitations. While the model detected unusual patterns associated with the obscured visual cues, it did not accurately capture and interpret the partially or fully obscured child. This limitation is noteworthy as it highlights a scenario where even humans might potentially miss. One possible reason is the current capacity of the visual encoders used in the MLLM.
While increasing encoder size could improve detection, it would also result in a larger model and slower real-time performance. A more efficient approach could be to use a lightweight visual encoder combined with a compact language model. By integrating temporal information from multiple frames, the model could better track subtle changes and movements, improving hazard detection without significantly increasing computational load.
VII Discussion
VII-A Domain Transfer
From the labeled response presented in Table II, proprietary LLM models demonstrate the ability to provide accurate and relevant answers to complex, domain-specific questions related to AD perception and SOTIF risks, even without prior specialized training. This highlights their effectiveness in zero-shot and one-shot scenarios, where little or no task-specific data is available.
Moreover, as seen in Tables V and VI-B, SFT processes enable baseline models to incorporate domain-specific knowledge. These fine-tuned models perform reasonably well when answering questions about SOTIF challenges for vehicles operating in both Waterloo, Canada, and China. This indicates that the models can effectively transfer learned knowledge across different environmental contexts, adapting to diverse traffic scenarios, weather conditions, and regulatory landscapes.
@@ VII-A Domain Transfer
VII-B Hallucinations
During the dataset generation process, we encountered cases of hallucinations in the labeled responses obtained before the validation agent. For instance, when asked whether a specific object was present in an image, half of the ten generated responses indicated “yes,” while the other half said “no.” This inconsistency highlights a common issue with large vision-language models (LVLMs), known as object hallucination [50].
After applying post-processing with LVLMs for validation, these hallucinations, particularly fact-conflicting hallucinations [50], can still appear in the dataset. Fine-tuned models may still inherit these hallucinations from their base LLM models.
Addressing hallucinations is a complex and open problem, especially in applications where responses must be both precise and trustworthy. Recent advancements, such as Fine-grained Direct Preference Optimization (FDPO) [51], Visual Contrastive Decoding (VCD) [52] and methods like HallE-Switch [53] have shown promise in reducing hallucination rates by 40-55%. Despite these improvements, hallucinations remain a persistent challenge. In the context of SOTIF, several strategies can be employed to mitigate these issues. These include integrating rule-based systems, incorporating human-in-the-loop validation, and developing hybrid models that combine statistical learning with formal verification methods.
VII-C Future of LLM in SOTIF
As current research on the application of LLMs in the SOTIF domain remains limited, this section lists potential future uses of both text-based and multimodal LLMs.
@@ VII-C Future of LLM in SOTIF
First, a universal description system for SOTIF scenarios can be developed to provide standardized descriptions of SOTIF scenarios. Building on this, a custom-trained or fine-tuned text-based LLM can be integrated to generate JSON-based representations for scenario detection, communication, and risk identification.
Text-based and multimodal LLMs could also be employed to efficiently generate SOTIF test cases, following systematic and structured testing plans. By leveraging LLM capabilities, these test plans can be progressively refined so that failure in one test case informs subsequent testing, ultimately reducing the total number of tests required.
Another promising application is the generation of highly targeted test cases. Given a specific SOTIF risk, LLMs can focus on creating scenarios that probe edge cases and rare, uncommon conditions, ensuring that the most challenging and potentially dangerous situations are thoroughly evaluated.
@@
Multimodal LLMs can assist in evaluating rule-based or learning-based SOTIF mechanisms. By providing advanced scene understanding and cognitive abilities, MLLMs can refine manually defined SOTIF evaluation algorithms and crosscheck the results from conventional systems, ensuring that SOTIF risks are properly assessed from multiple angles. Additionally, MLLMs can select specific scenarios of interest based on input prompts.
@@
The potential application of multimodal LLMs lies in integrating data from multiple sensor types, such as lidar, radar, and bird-eye-view (BEV) images. By encoding these inputs into feature embeddings and mapping them alongside textual prompts, LLMs could directly assess SOTIF-related risks at the feature level. This would provide a more detailed understanding of risks, eliminating the reliance on processed detection results and potentially identifying hazards earlier in the pipeline.
VIII Conclusion and Future Work
@@
Our research demonstrates the feasibility of applying SFT techniques to MLLMs using the DriveSOTIF dataset to enhance perception-related SOTIF risk detection in autonomous driving systems. By introducing a small domain-specific dataset intended for model fine-tuning, we were able to evaluate strengths and limitations of MLLMs in perception-related SOTIF problems. Benchmark results indicate that our fine-tuned MLLMs outperform baseline models both qualitatively and quantitatively.
@@
These improvements lead to enhanced risk identification and situational awareness, even under challenging conditions such as adverse weather and occlusion-heavy scenarios, as observed in data collected from both China and Canada. Real-world performance testing further confirms that the deployed models achieve an average inference time of 0.82 seconds per image, making them suitable for low- to moderate-speed applications like urban navigation, automated parking, and traffic jam assistance.
@@
More importantly, by integrating fine-tuned MLLMs with LLM-based decision-making frameworks, autonomous vehicles can incorporate perception risks into real-time operations, ultimately improving road safety. Moreover, the ability of MLLMs to interpret complex scenarios through natural language reasoning paves the way for more transparent and interpretable safety assessments. Lastly, the on-device or MEC-based deployment of these lightweight models can lead to a privacy-preserving, onboard-cloud hybrid LLM deployment for next-generation autonomous vehicles.
For future work, we plan to expand the DriveSOTIF dataset by incorporating additional modalities such as BEV images, LiDAR point clouds, and radar data, which will further enhance the spatial-temporal reasoning capabilities of MLLMs. Our study also highlights several challenges that warrant further investigation, including effectively managing occlusion-heavy scenarios, scaling the technology for high-speed and safety-critical applications, and reducing the dependence on human review while ensuring data quality and minimizing hallucinations. Addressing these challenges will require collaborative efforts at the intersection of natural language processing and autonomous driving research.
@@ Acknowledgment
The authors would like to acknowledge the financial support of the Natural Sciences and Engineering Research Council of Canada (NSERC) and the grant provided by OpenAI.

### v2

@@ V-B Real World Data Evaluation
For real-world evaluation, as shown in Table V-B, we evaluated the fine-tuned models using data collected in Waterloo, Ontario, Canada, along with proprietary dash-cam footage from China. This evaluation is conducted through a ROS code that queries the MLLM API with random questions from a list, along with image data from collected data in the form of ROSBag or video.
@@
In detail, we used image and video data from the University of Waterloo’s WATonoBus project [45] and the Canadian Adverse Driving Conditions (CADC) dataset [46]. The WATonoBus is an autonomous shuttle operating along a 2.7-kilometer, two-lane University Ring Road that encircles the campus. The CADC dataset was gathered by an autonomous vehicle platform navigating various areas around Waterloo.
@@
TABLE IX: Results of Real-World Evaluation
@@
A: The risks include potential misidentification of road boundaries due to snow, reduced contrast between the road and its surroundings, and the presence of workers and equipment which could be obscured by snow or poorly illuminated. A: The perception-related SOTIF risk in this image is caused by heavy snowfall, which can reduce visibility and obscure road markings, signs, and other vehicles.
@@
V-C Results
As shown in Table V-B, in our experiments, fine-tuned MLLMs performed effectively in challenging environments, yet it also highlighted critical areas for enhancement. Notably, in Waterloo’s snowy conditions, the MLLM successfully managed complex scenarios with obscured road markers and temporary obstacles like snowplows and maintenance crews. This performance highlights the model’s robust feature recognition capabilities under adverse weather conditions, proving its potential in dynamic real-world applications.
@@ V-C Results
Conversely, the MLLM’s performance in detecting subtle hazards, such as the partially visible child in a box during the China experiments, revealed some limitations. While the model detected unusual patterns associated with the obscured visual cues, it did not accurately capture and interpret the partially or fully obscured child. This highlights a scenario that is challenging even for human drivers. One possible reason is the current capacity of the visual encoders used in the MLLM.
VI Discussion
VI-A Implications of Experimental Results
The experimental findings highlight three key capabilities: accurate risk detection and cognition under adverse condition, causal reasoning linking environmental factors to safety outcomes, and actionable mitigation strategies that address specific SOTIF vulnerabilities. With only 2 GB of GPU memory, the fine-tuned InternVL3-1B model can be deployed to vehicles and replace the API-based vision model for local (near) real-time visual risk assessment. This eliminates privacy concerns, cybersecurity risks, and reliance on internet access.
Pairing this model with a decision-making framework like DriveLLM[45] enables autonomous systems to incorporate SOTIF risks into real-time decision-making. This integration improves safety in complex environments and potentially prevents fatal accidents like the child-hidden-in-the-box scenario, where conventional perception systems may fail to anticipate hidden risks.
Most importantly, deploying a fine-tuned model that supports continuous improvement significantly enhances adaptability. By continuously updating the dataset and applying techniques like SFT and reinforcement learning from human feedback (RLHF) [47], the MLLM can progressively refine its performance and extend to other tasks such as scenario understanding and ODD assessment.
VI-B Domain Transfer
From the labeled response presented in Table III, proprietary LLM models demonstrate the ability to provide accurate and relevant answers to complex, domain-specific questions related to AD perception and SOTIF risks, even without prior specialized training. This highlights their effectiveness in zero-shot and one-shot scenarios, where little or no task-specific data is available.
Moreover, as seen in Tables V and V-B, SFT processes enable baseline models to incorporate domain-specific knowledge. These fine-tuned models perform reasonably well when answering questions about SOTIF challenges for vehicles operating in both Waterloo, Canada, and China. This indicates that the models can effectively transfer learned knowledge across different environmental contexts, adapting to diverse traffic scenarios, weather conditions, and regulatory landscapes.
@@ VI-B Domain Transfer
VI-C Hallucination
Hallucination in LLMs refers to the generation of plausible yet factually incorrect or inconsistent responses. It represents a critical challenge for safety-critical applications like autonomous driving. Hallucination can occur at various stages, including pre-training, post-training, SFT, and inference[48]. Within this research, we focus specifically on hallucination suppression during the SFT stage.
During dataset generation, we observed instances of object hallucination, where VLMs would inconsistently report the presence of objects across multiple queries. For example, when asked whether a specific object was present in an image, one out of three attempts indicated ”yes,” while the remaining attempts responded ”no.” This inconsistency exemplifies fact-conflicting hallucination, a critical reliability issue in VLMs where models generate contradictory responses to identical queries [48].
To address this challenge, we implemented a multi-stage hallucination suppression strategy during the SFT process. Our approach includes: (1) a post-processing validation step using an independent LLM to cross-verify generated responses against input images, (2) consistency filtering that removes outputs failing validation checks, and (3) iterative correction of inconsistent annotations. This targeted intervention significantly reduces hallucinated content entering the fine-tuning pipeline, ensuring our dataset maintains grounding in factual, image-based annotations.
However, inference-time hallucination remains a persistent challenge, particularly due to inherent knowledge gaps and biases in pre-trained model backbones. Existing mitigation strategies present distinct trade-offs for real-time applications. Retrieval-augmented generation [49], while effective at grounding responses in factual content, introduces additional inference latency that may be incompatible with the sub-second response requirements of autonomous driving systems. Human feedback mechanisms [50], though widely adopted in industry and demonstrably effective at improving model reliability, face scalability constraints when applied to high-frequency inference scenarios typical in autonomous vehicles. At the moment, one feasible direction for addressing inference-time hallucination is grounding model response through a human-in-loop semi-auto feedback mechanism. Other methods, like latent space steering [51], can also be explored to reduce hallucination while maintaining real-time performance.
VI-D Future of LLM in SOTIF
@@
As current research on LLMs in SOTIF applications remains limited, this section outlines potential future uses of both text-based and multimodal LLMs. First, a universal description system for SOTIF scenarios can be developed to provide standardized descriptions. Building on this foundation, custom-trained LLMs can generate JSON-based representations for scenario detection, communication, and risk identification. This standardization would make it easier to share and compare SOTIF scenarios across different research groups and industry applications.
These standardized descriptions would then support better testing approaches. LLMs can efficiently generate SOTIF test cases following systematic testing plans. By learning from test failures, these systems can improve subsequent tests, reducing the total number of tests needed. LLMs can also create highly targeted test cases for specific SOTIF risks, focusing on edge cases and rare conditions that are most challenging and potentially dangerous.
@@
The testing capabilities naturally extend to evaluation tasks. Multimodal LLMs can assist in evaluating existing rule-based or learning-based SOTIF methods. By providing advanced scene understanding, MLLMs can refine manually defined evaluation algorithms and cross-check results from conventional systems. This ensures SOTIF risks are assessed from multiple angles. MLLMs can also select specific scenarios of interest based on input prompts.
@@
Finally, multimodal LLMs can integrate data from multiple sensor types, such as LiDAR, radar, and bird’s-eye-view images. By encoding these inputs into feature embeddings alongside textual prompts, LLMs could directly assess SOTIF risks at the feature level. This approach would provide more detailed risk understanding, reduce reliance on processed detection results, and potentially identify hazards earlier in the perception pipeline.
VII Conclusion and Future Work
@@
Our research demonstrates that SFT techniques applied to MLLMs using the DriveSOTIF dataset can enhance perception-related SOTIF risk detection in autonomous driving systems. By introducing a small domain-specific dataset intended for model fine-tuning, we were able to evaluate strengths and limitations of MLLMs in perception-related SOTIF problems. Benchmark results indicate that our fine-tuned MLLMs outperform baseline models both qualitatively and quantitatively.
@@
These improvements enhance risk identification and situational awareness, even under challenging conditions such as adverse weather and occlusion-heavy scenarios. This is confirmed by experiments on data collected from both China and Canada. Real-world performance testing further confirms that the deployed models achieve an average inference time of 0.59 second per image, making them suitable for low- to moderate-speed applications like urban navigation, automated parking, and traffic jam assistance.
@@
More importantly, by integrating fine-tuned MLLMs with LLM-based decision-making frameworks, autonomous vehicles can incorporate perception risks into real-time operations, ultimately improving road safety. Moreover, the ability of MLLMs to interpret complex scenarios through natural language reasoning establishes concrete foundations for transparent safety assessment through natural language explanations of risk factors and mitigation strategies. Lastly, the on-device and MEC-based deployment of these lightweight models can lead to a privacy-preserving, onboard-cloud hybrid LLM deployment for next-generation autonomous vehicles.
For future work, we plan to improve real-time performance for high-speed safety-critical applications, and minimize hallucinations during inference stage. This will require interdisciplinary collaboration spanning natural language processing and autonomous driving to realize the full potential of MLLMs in safer autonomous driving systems.
@@ Acknowledgment
The authors would like to acknowledge the financial support of the Natural Sciences and Engineering Research Council of Canada (NSERC) and Vector Institute. We also acknowledge the research grant provided by OpenAI.